OpenStudy (anonymous):

will give medal Find the value of the expression 2x3 + 3y2 − 17 when x = 3 and y = 4. (5 points) 26 85 943 1,927

OpenStudy (anonymous):

Okay so you need to substitute. You know how?

OpenStudy (anonymous):

You just insert the x=? and the y=? in for the x's and y's in the equation.

OpenStudy (anonymous):

2x3 +3y2 − 17 =2 *3*3 + 3*4*2-17 =25

OpenStudy (anonymous):

25 isn't even an answer up there

OpenStudy (welshfella):

2x3 + 3y2 − 17 x = 3 y = 4 so we have 2(3)^3 + 3(4)^2 - 17 = 54 + 48 - 17

OpenStudy (welshfella):

i've assumed that the numbers after the x an y are exponents

OpenStudy (anonymous):

So the answer would be 85

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

substitute the 3 for x and 4 for y and do the math it would come out to 85

OpenStudy (anonymous):

I see now that makes total sense thank you so very much

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

np

OpenStudy (anonymous):

does any body know how to do this one If the length of side a is 12 centimeters, m∠B = 36°, and m∠C = 75°, what is the length of side b? Round your response to two decimal places.